# Felix Arnold Huawei, Switzerland    Maxence Bouvier Huawei, Switzerland    Ryan Amaudruz Huawei, Switzerland    Renzo Andri Huawei, Switzerland    Lukas Cavigelli Huawei, Switzerland

### v2

@@ IV Results
In the following, we deploy the EPFL Combinational Benchmark [7]. Starting with the original Verilog implementation, we apply our method targeting MIG nodes minimization, and Table II presents the resulting minimal circuit. We demonstrate an improvement of $20.94\%$ in MIG minimization on the max.v circuit. As the duration of each step grows dramatically with the design size (from 5 ms to 21 min for some recipes), we have not yet applied the full method on larger designs. And-Inverter Graph (AIG)node counts of the MIG optimized circuits are added for reference, but have not been optimized for. The synthesized designs are available on GitHub (link).
@@ V Conclusion
The proposed method shows a 14$\times$ synthesis speedup and 20.94% better MIG minimization through the integration of predictive models and iterative selection strategies. Hyperparameter optimization remains crucial, as the best configuration varies among circuits. Since a considerable number of steps are required to achieve the best QoR, minimization of large designs remains challenging. We believe that further QoR improvement can be achieved with reinforcement learning methods. As the PrM, PoM, and IISM are originally independent of the MT software, our method could be deployed with other tools.

### v3

# The Art of Beating the Odds with

@@ IV Results
In the following, we deploy the EPFL Combinational Benchmark [7]. Starting with the original Verilog implementation, we apply our method targeting MIG nodes minimization, and Table II presents the resulting minimal circuit. We demonstrate an improvement of up to $20.94\%$ in MIG minimization on the max.v circuit. As the duration of each step grows dramatically with the design size (from 5 ms to 21 min for some recipes), we have not yet applied the full method on larger designs. And-Inverter Graph (AIG)node counts of the MIG optimized circuits are added for reference, but have not been optimized for. The synthesized designs are available on GitHub.
@@ V Conclusion
The proposed method shows a 14$\times$ synthesis speedup and up to 20.94% better MIG minimization through the integration of predictive models and iterative selection strategies. Hyperparameter optimization remains crucial, as the best configuration varies among circuits. Since a considerable number of steps are required to achieve the best QoR, minimization of large designs remains challenging. We believe that further QoR improvement can be achieved with reinforcement learning methods. As the PrM, PoM, and IISM are originally independent of the MT software, our method could be deployed with other tools.